# Solvable Tuple Patterns and Their Applications to Program Verification

### v1

@@ A.1.7 Alternative Proof of Minimality
Let $t$ be a solvable tuple pattern, and $\Theta$ be $\left(\begin{array}[]{ccc}x_{1}&\cdots&x_{k}\\ \alpha_{1}&\cdots&\alpha_{k}\\ \beta_{1}&\cdots&\beta_{k}\\ \end{array}\right)$, where $\alpha_{i},\beta_{i}$ are prefix codewords for $x_{i}$ as defined above. If $\Theta^{\prime}t=M^{t}$, then $\Theta^{\prime}(x_{i})=\Theta(x_{i})$ for every $x_{i}\in\mathbf{Vars}(t)$.
@@
Let $t=(p_{1},\ldots,p_{n})$ be a solvable tuple pattern such that $|p_{1}\cdots p_{n}|=m$. We show $M=M^{t}$ satisfies the required properties. By the definition of $M^{t}$, $\mathit{size}(M^{t})=O((m+n)\log n)$. By $M^{t}\models_{s}t$ and Theorem 2.4, we have $(\widetilde{x},[M^{t}/\widetilde{x}])\longrightarrow^{*}(t,\Theta)$ for some $\Theta$. For (ii), it remains to show $(t,\Theta)\;\mathbin{\ \not\!\!\longrightarrow}$. By Theorem 2.1, we have $\Theta t=M^{t}$. By Lemma A.19, $\Theta$ must be:
$\left(\begin{array}[]{ccc}x_{1}&\cdots&x_{k}\\ \alpha_{1}&\cdots&\alpha_{k}\\ \beta_{1}&\cdots&\beta_{k}\\ \end{array}\right).$
@@
Because $\alpha_{i}$’s and $\beta_{i}$’s are prefix codewords where $a$ and $b$ are swapped, no rule is applicable to reduce $(t,\Theta)$; hence we have $(t,\Theta)\;\mathbin{\ \not\!\!\longrightarrow}$ as required. To show (iii), suppose $(\widetilde{x},[M^{t}/\widetilde{x}])\longrightarrow^{*}(t^{\prime},\Theta^{\prime})\;\mathbin{\ \not\!\!\longrightarrow}$. By Theorem 2.1, $t^{\prime}\models M^{t}$. By Lemma A.18, we have $\mathcal{L}(t)\subseteq\mathcal{L}(t^{\prime})$. By Theorem 2.5, we have $\mathcal{L}(t)=\mathcal{L}(t^{\prime})$ as required. Obviously, the size of the canonical data $M^{t}$ is $O((m+n)\log n)$, and it can be computed also in time $O((m+n)\log n)$. ∎
@@
(1): By Lemma A.5, if $t\leadsto t^{\prime}$, then $t$ is solvable if and only if $t^{\prime}$ is solvable. By Lemma A.6, the length of a reduction sequence $t\leadsto t_{1}\leadsto\cdots\leadsto t_{n}$ is bounded by #t\#{t}. Thus, to check whether $t$ is solvable, it suffices to repeatedly reduce $t$ (in a non-deterministic manner) until no reduction is applicable, and check weather the resulting tuple pattern is of the form $(x_{1},\ldots,x_{k})$. As each step can be performed in polynomial time, the algorithm runs in polynomial time.
(3): Based on Lemma A.4, we can use the following algorithm:
1:function decide_inc($t_{1},t_{2}$)
2:  if $t_{2}$ is of the form $(x_{1},\ldots,x_{n})$ (where $x_{1},\ldots,x_{n}$ are mutually different) then
3:   return TRUE;
4:  end if
5:  pick $t_{2}^{\prime}$ such that $t_{2}\leadsto t_{2}^{\prime}$;
6:  if $t_{1}^{\prime}:=\langle t_{2}\leadsto t_{2}^{\prime}\rangle(t_{1})$ is well defined then
7:   return decide_inc($t_{1}^{\prime},t_{2}^{\prime}$);
8:  else
9:   return FALSE;
10:  end if
11:end function
@@
Here, we assume $|t_{1}|=|t_{2}|$; otherwise, we can immediately conclude that $\mathcal{L}(t_{1})\not\subseteq\mathcal{L}(t_{2})$. The algorithm repeatedly reduces $t_{2}$ and checks whether the reduction of $t_{2}$ can be “simulated by” the corresponding reduction of $t_{1}$. On line 3, we can immediately conclude that $t_{1}\subseteq t_{2}$ holds, as $\mathcal{L}(t_{2})=\Sigma^{*}\times\cdots\times\Sigma^{*}$. On line 5, we can always pick $t_{2}^{\prime}$ by the assumption that $t_{2}$ is solvable. By Lemma A.4, if $\langle t_{2}\leadsto t_{2}^{\prime}\rangle(t_{1})$ is well defined, then it suffices to check whether $\mathcal{L}(t_{1}^{\prime})\subseteq\mathcal{L}(t_{2}^{\prime})$ (line 7); otherwise, we can immediately conclude that $\mathcal{L}(t_{1})\not\subseteq\mathcal{L}(t_{2})$ (line 9).
@@
Each operation can be clearly performed in time polynomial in the size of the initial inputs $t_{1}$ and $t_{2}$. Since ${t_{2}}>\#{t_{2}^{\prime}}\geq 0$, the depth of the recursive call is bounded by $t_{2}$}. Thus, the algorithm runs in polynomial time.
@@
A.2 Proofs for the Extensions in Section 2.3
A.2.1 Proofs for the extension in Section 2.3.1.
Due to the symmetry between prefix and postfix, all the properties in Table 3 except Lemma A.5 follow immediately. Thus it suffices to re-prove Lemma A.5.
We prepare the following lemma.
Lemma A.20.
[…]
Proof.

### v2

@@ A.1.3 Correspondence between $\longrightarrow$ and $\leadsto$
Let $t$ be a solvable tuple pattern, and $\Theta$ be $\left(\begin{array}[]{ccc}x_{1}&\cdots&x_{k}\\ \alpha_{1}&\cdots&\alpha_{k}\\ \beta_{1}&\cdots&\beta_{k}\\ \end{array}\right)$, where $\alpha_{i},\beta_{i}$ are the codewords for $x_{i}$ as defined above. If $\Theta^{\prime}t=M^{t}$, then $\Theta^{\prime}(x_{i})=\Theta(x_{i})$ for every $x_{i}\in\mathbf{Vars}(t)$.
@@
Let $t=(p_{1},\ldots,p_{n})$ be a solvable tuple pattern such that $|p_{1}\cdots p_{n}|=m$. We show $M=M^{t}$ satisfies the required properties. By the definition of $M^{t}$, $\mathit{size}(M^{t})=O((m+n)\log n)$. Suppose $M\subseteq M^{\prime}\subseteq\mathcal{L}(t)$. We can assume that the first two rows of $M^{\prime}$ (as a matrix) coincide with $M$. By $M^{\prime}\models_{s}t$ and Theorem 3.2, we have $(\widetilde{x},[M^{t}/\widetilde{x}])\longrightarrow^{*}(t,\Theta)$ for some $\Theta$. For (ii), it remains to show $(t,\Theta)\;\mathbin{\ \not\!\!\longrightarrow}$. By Theorem 2.1, we have $\Theta t=M^{\prime}$. By Lemma A.15, $\Theta$ must be:
$\left(\begin{array}[]{ccc}x_{1}&\cdots&x_{k}\\ \alpha_{1}&\cdots&\alpha_{k}\\ \beta_{1}&\cdots&\beta_{k}\\ \cdots&\cdots&\cdots\end{array}\right).$
@@
Because there exist no $i$ and $j$ such that $\alpha_{i}$ is a prefix or suffix of $\alpha_{j}$, the rules R-Prefix and R-Suffix are inapplicable. Since $a$ and $b$ are swapped between $\alpha_{i}$’s and $\beta_{i}$’s, the rules R-CPrefix and R-CSuffix are also inapplicable. Thus, we have $(t,\Theta)\;\mathbin{\ \not\!\!\longrightarrow}$ as required. To show (iii), suppose $(\widetilde{x},[M^{\prime}/\widetilde{x}])\longrightarrow^{*}(t^{\prime},\Theta^{\prime})\;\mathbin{\ \not\!\!\longrightarrow}$. By Theorem 2.1, $t^{\prime}\models M^{\prime}$. By Lemma A.14, we have $\mathcal{L}(t)\subseteq\mathcal{L}(t^{\prime})$. By Theorem 3.4, we have $\mathcal{L}(t)=\mathcal{L}(t^{\prime})$ as required. Obviously, the size of the characteristic data $M^{t}$ is $O((m+n)\log n)$, and it can be computed also in time $O((m+n)\log n)$. ∎
A.1.8 Proofs for Section 3.3
Lemma A.16.
Suppose $M^{\prime}\subseteq M$ and $(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{*}(t,\Theta)$. Then there exist $t^{\prime}$ and $\Theta^{\prime}$ such that $(\widetilde{x},[M^{\prime}/\widetilde{x}])\longrightarrow^{*}(t^{\prime},\Theta^{\prime})$ and $t^{\prime}=[\widetilde{\epsilon}/\widetilde{y}]t$ for some variables $\widetilde{y}$.
Proof.
This follows by induction on the length $n$ of the reduction sequence $(\widetilde{x},M)\longrightarrow^{*}(t,\Theta)$. If $n=0$, then $(\widetilde{x},[M/\widetilde{x}])=(t,\Theta)\;\mathbin{\ \not\!\!\longrightarrow}$. Thus, the required result holds for the empty sequence of variables $\widetilde{y}$.
If $n>0$, then we have $(\widetilde{x},[M/\widetilde{x}])\longrightarrow(t_{1},[M_{1}/\widetilde{y}])\longrightarrow^{n-1}(t,\Theta)$. By Lemma A.1, we have $(\widetilde{y},[M_{1}/\widetilde{y}])\longrightarrow^{n-1}(t_{2},\Theta)$ with $t=[t_{2}/\widetilde{y}]t_{1}$.
We perform case analysis on the rule used for deriving $(\widetilde{x},M)\longrightarrow(t_{1},\Theta_{1})$. Since the other cases are similar or easier, we discuss only the cases for R-Epsilon and R-Prefix.
  • Case R-Epsilon: In this case, we have:
    $M[*][i]=\widetilde{\epsilon}$
    $(\widetilde{y})=(x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{k})$
    $t_{1}=(x_{1},\ldots,x_{i-1},\epsilon,x_{i+1},\ldots,x_{k})$
    $M_{1}=M\mathbin{\uparrow_{i}}.$
[…]
  • Case R-Prefix: In this case, We have:
    $M[*][i]\neq\widetilde{\epsilon}$
    $M[*][j]=M[*][i]\cdot\widetilde{s}$
    $(\widetilde{y})=(x_{1},\ldots,x_{j-1},x_{j}^{\prime},x_{j+1},\ldots,x_{k})$
    $t_{1}=(x_{1},\ldots,x_{j-1},x_{i}x_{j}^{\prime},x_{j+1},\ldots,x_{k})$
    $M_{1}=M\{j\mapsto\widetilde{s}\}$.
[…]
For data $M$ such that $\mathit{dim}({M})=n$, we define $\mathit{CSTPs}(M)$ as $\mathit{normalize}(C_{0},\ldots,C_{n})$ where:
$C_{i}=\{t\mid(\widetilde{x},[M/\widetilde{x}]\longrightarrow^{*}(t,\Theta)\mbox{ and }|\mathbf{Vars}(t)|=i\}$
$\mathit{normalize}(C_{0},\ldots,C_{n})=(C^{\prime}_{0},\ldots,C^{\prime}_{n})$
[…]
For example, for $M=\left(\begin{array}[]{rrr}a&ab&abab\\ bb&b&bb\end{array}\right)$, $\mathit{CSTPs}(M)$ is:
$(\emptyset,\emptyset,\{(x,y,yy)\},\{(x,y,z),(x,y,xz),(x,y,yz),(x,y,zy)\}).$
(Note that we also have $((x,y,z),[M/(x,y,z)])\longrightarrow^{*}((x,y,yyz),\Theta)$, but $(x,y,yyz)$ is subsumed by $(x,y,yy)$, as $\epsilon/z](x,y,yyz)=(x,y,yy)$.)
We claim:
Lemma A.17.
Suppose $\dim(M)=n$. If $\mathit{CSTPs}(M)=(C_{0},\ldots,C_{n})$, then $\mathcal{L}(\mathtt{CTPinf}(M))=\mathcal{L}(\bigwedge_{t\in C_{0}\cup\cdots\cup C_{n}}t)$.
Proof.
To show $\mathcal{L}(\mathtt{CTPinf}(M))\supseteq\mathcal{L}(\bigwedge_{t\in C_{0}\cup\cdots\cup C_{n}}t)$, suppose $t$ is a conjunct of $\mathtt{CTPinf}(M)$. Since $t$ contains at most $n$ variables, by the definition of $\mathit{CSTPs}(M)$, there exists $t^{\prime}\in C_{j}$ such that $t^{\prime}=[\widetilde{\epsilon}/\widetilde{y}]t$. Thus, $\mathcal{L}(t)\supseteq\mathcal{L}(t^{\prime})\supseteq\mathcal{L}(\bigwedge_{t\in C_{0}\cup\cdots\cup C_{n}}t)$. Therefore, we have $\mathcal{L}(\mathtt{CTPinf}(M))\supseteq\mathcal{L}(\bigwedge_{t\in C_{0}\cup\cdots\cup C_{n}}t)$, as required.
For the converse, suppose $t\in C_{i}$. Then, we have $(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{*}(t,\Theta)$ for some $\Theta$. Let $t^{\prime}$ be a tuple pattern such that $(t,\Theta)\longrightarrow^{*}(t^{\prime},\Theta^{\prime})\;\mathbin{\ \not\!\!\longrightarrow}$. Then, $t^{\prime}$ is a conjunct of $\mathtt{CTPinf}(M)$, and therefore, we have $\mathcal{L}(\mathtt{CTPinf}(M)\subseteq\mathcal{L}(t^{\prime})\subseteq\mathcal{L}(t)$. Thus, we have $\mathcal{L}(\mathtt{CTPinf}(M))\subseteq\mathcal{L}(\bigwedge_{t\in C_{0}\cup\cdots\cup C_{n}}t)$, as required. ∎
We define the order $(C_{0},\ldots,C_{n})<(C^{\prime}_{0},\ldots,C^{\prime}_{n})$ by: $(C_{0},\ldots,C_{n})<(C^{\prime}_{0},\ldots,C^{\prime}_{n})$ iff $\exists i.(C_{i}\subsetneq C^{\prime}_{i}\land\forall j<i.C_{j}=C^{\prime}_{j})$. Note that << is well-founded.
Lemma A.18.
Suppose $M_{0},M_{1},M_{2},\ldots$ is an infinite, strictly decreasing sequence of non-empty data, i.e., $\emptyset\subsetneq M_{0}\subsetneq M_{1}\subsetneq M_{2}\subsetneq\cdots$. Then, there exists $j$ such that for all $i\geq j$, $\mathcal{L}(\mathtt{CTPinf}(M_{i}))=\mathcal{L}(\mathtt{CTPinf}(M_{j}))$.
Proof.
By Lemma A.16, we have:
$\mathit{CSTPs}(M_{0})\geq\mathit{CSTPs}(M_{1})\geq\mathit{CSTPs}(M_{2})\geq\cdots$
Since << is well-founded, there exists $j$ such that for all $i\geq j$, $\mathit{CSTPs}(M_{i})=\mathtt{CTPinf}(M_{j})$. By Lemma A.17, we have $\mathcal{L}(\mathtt{CTPinf}(M_{i}))=\mathcal{L}(\mathtt{CTPinf}(M_{j}))$ for all $i\geq j$. ∎
@@
Suppose $t$ is an STP, and $\mathcal{L}(t)=\{\widetilde{s}_{i}\mid i\in\omega\}$ with $M_{i}=\{\widetilde{s}_{j}\mid 0\leq j\leq i\}$. By Theorem 3.5, there exists $M$ such that $\mathtt{TPinf}(M^{\prime})$ always returns an STP equivalent to $t$ for any $M^{\prime}\supseteq M$, and by the soundness of $\mathtt{TPinf}$ (Theorem 2.1), we have $M\subseteq\mathcal{L}(t)$. Since $\mathcal{L}(t)=\{\widetilde{s}_{i}\mid i\in\omega\}$, there exists $j$ such that $M\subseteq M_{j}$. Therefore, for every $k\geq j$, $\mathtt{TPinf}(M_{k})$ always returns an STP equivalent to $t$.
To show the case for CSTPs, suppose $\gamma$ is a CSTP and $\mathcal{L}(\gamma)=\{\widetilde{s}_{i}\mid i\in\omega\}$ with $M_{i}=\{\widetilde{s}_{j}\mid 0\leq j\leq i\}$. By Lemma A.18, there exist $\gamma^{\prime}$ and $j$ such that for all $i\geq j$, $\mathcal{L}(\mathtt{CTPinf}(M_{i}))=\mathcal{L}(\gamma^{\prime})$. By the soundness of $\mathtt{TPinf}$ (Theorem 2.1), we have $\mathcal{L}(\gamma)=\bigcup_{i\in\omega}M_{i}\subseteq\mathcal{L}(\gamma^{\prime})$. By Theorem 3.3, we also have $\mathcal{L}(\gamma)\supseteq\mathcal{L}(\gamma^{\prime})$. Therefore, we have $\mathcal{L}(\gamma)=\mathcal{L}(\gamma^{\prime})$, as required. ∎
Proof of Theorem 3.7.
We first show the termination of the procedure by contradiction. Suppose the procedure does not terminate. Then, there exist infinite sequences $\widetilde{s}_{0},\widetilde{s}_{1},\widetilde{s}_{2},\ldots$ and $\gamma_{0},\gamma_{1},\gamma_{2},\ldots$ such that:
$\gamma_{0}=(a,\widetilde{\epsilon})\land(\epsilon,\widetilde{\epsilon})\qquad M_{j}=\{\widetilde{s}_{i}\mid i<j\}.$
$\mathtt{CTPinf}(M_{j})=\gamma_{j}\mbox{ for $j\geq 1$}$
$g_{\mathcal{F}}(\gamma_{j})=Some(\widetilde{s}_{j})\mbox{ for $j\geq 0$}.$
These $\gamma_{j}$, $\widetilde{s}_{j}$, and $M_{j}$ are the values of $\gamma$, $\widetilde{s}$, and $M$ at the $j$-th iteration of the loop in the procedure. By the last condition and the assumption on $g_{\mathcal{F}}$, we have $\widetilde{s}_{j}\in\mathcal{F}(\mathcal{L}(\gamma_{j}))\setminus\mathcal{L}(\gamma_{j})$. By the soundness of $\mathtt{CTPinf}$ (Theorem 2.1), we have $M_{j}\subseteq\mathcal{L}(\gamma_{j})$. Therefore, $\widetilde{s}_{j}\not\in M_{j}$, and hence, $M_{0}=\emptyset\subsetneq M_{1}\subsetneq M_{2}\subsetneq M_{3}\subsetneq\cdots$ is a strictly increasing sequence. By Lemma A.18, there exist $k$ and $\gamma$ such that $\mathcal{L}(\gamma_{j})=\mathcal{L}(\gamma)$ for all $j\geq k$. This however, contradicts $\widetilde{s}_{j}\in\mathcal{F}(\mathcal{L}(\gamma_{j}))\setminus\mathcal{L}(\gamma_{j})$ and $\widetilde{s}_{j}\in M_{j+1}\subseteq\mathcal{L}(\gamma_{j+1})$.
Next, we show that the procedure returns the least CSTP $\gamma$ such that $\mathcal{L}(\gamma)\supseteq\mathcal{F}(\mathcal{L}(\gamma))$. Suppose $\mathcal{L}(\gamma^{\prime})\supseteq\mathcal{F}(\mathcal{L}(\gamma^{\prime}))$ for a CSTP $\gamma^{\prime}$. Let $\gamma_{j}$, $\widetilde{s}_{j}$, and $M_{j}$ be the values of $\gamma$, $\widetilde{s}$, and $M$ at the $j$-th iteration of the loop in the procedure, as given above. It suffices to show $\mathcal{L}(\gamma_{j})\subseteq\mathcal{L}(\gamma^{\prime})$ by induction on $j$. The case $j=0$ is obvious, as $\mathcal{L}(\gamma_{j})=\emptyset$. For the induction step, suppose $\mathcal{L}(\gamma_{j})\subseteq\mathcal{L}(\gamma^{\prime})$ and $g_{\mathcal{F}}(\gamma_{j})$ returns $Some(\widetilde{s}_{j})$. By the monotonicity of $\mathcal{F}$, we have
$\mathcal{F}(\mathcal{L}(\gamma_{j}))\subseteq\mathcal{F}(\mathcal{L}(\gamma^{\prime}))\subseteq\mathcal{L}(\gamma^{\prime}).$
Thus, $\widetilde{s}_{j}\in\mathcal{F}(\mathcal{L}(\gamma_{j}))\setminus\mathcal{L}(\gamma_{j})\subseteq\mathcal{L}(\gamma^{\prime})$. Therefore, we have
$M_{j+1}=M_{j}\cup\{\widetilde{s}_{j}\}\subseteq\mathcal{L}(\gamma_{j})\cup\{\widetilde{s}_{j}\}\subseteq\mathcal{L}(\gamma^{\prime}).$
By Theorem 3.3, we have $\gamma_{j+1}=\mathtt{CTPinf}(M_{j+1})\subseteq\mathcal{L}(\gamma^{\prime})$, as required. ∎
Proof of Theorem 3.8.
Let $\mathtt{enumerate}$ be the procedure defined as below, and suppose $\widetilde{s}_{0}\in\mathcal{F}(\emptyset)$. Then $\mathtt{enumerate}(\{\widetilde{s}_{0}\})$ outputs a superset of all the minimal $t$’s such that $\mathcal{L}(t)\supseteq\mathcal{F}(\mathcal{L}(t))$. We can then filter out non-minimal STPs by using the algorithm of Theorem 3.9 (3).
Function enumerate($M$):
 $U$ := $\{t\mid(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{*}(t,\Theta)\;\mathbin{\ \not\!\!\longrightarrow}\}$;
 while $U\neq\emptyset$ do
    $t$ := pick($U$); $U$ := $U\setminus\{t\}$;
[…]
       $U:=U\setminus\{t^{\prime}\mid\widetilde{s}\in\mathcal{L}(t^{\prime})\}$;
       enumerate ($M\cup\{\widetilde{s}\}$);
    else
       $\mathit{output}$($t$);
   end while
Note that $U$ in the above algorithm is a local variable, prepared for each recursive call.
We first show the termination of the algorithm. Suppose $\mathtt{enumerate}(\{\widetilde{s}_{0}\})$ does not terminate. Since $\{t\mid(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{*}(t,\Theta)\;\mathbin{\ \not\!\!\longrightarrow}\}$ is a finite set, there must be an infinite chain of recursive calls $\mathtt{enumerate}(M_{0}),\mathtt{enumerate}(M_{1}),\mathtt{enumerate}(M_{2}),\ldots$ where $M_{0}=\{\widetilde{s}_{0}\}$ and $M_{i}=M_{i-1}\cup\{\widetilde{s}_{i}\}$ for $i>0$. Let $U_{i}$ be the value of $U$ initialized at the beginning of the call $\mathtt{enumerate}(M_{i})$, and $t_{i}$ be the value of $t$ when $\mathtt{enumerate}(M_{i+1})$ is called. Thus, $M_{i}\subseteq\mathcal{L}(t_{i})$ and $\widetilde{s}_{i+1}\not\in\mathcal{L}(t_{i})$. By Theorem 3.6, there exists $k$ and $t^{\prime}$ such that $\mathcal{L}(t_{i})=\mathcal{L}(t^{\prime})$ for all $i\geq k$. But this contradicts the conditions $\widetilde{s}_{i}\in M_{i}\subseteq\mathcal{L}(t_{i})$ but $\widetilde{s}_{i}\not\in\mathcal{L}(t_{i-1})$.
Next, suppose that $t$ is a minimal STP such that $\mathcal{L}(t)\supseteq\mathcal{F}(\mathcal{L}(t))$. It suffices to show that for any $M$ such that $M\subseteq\mathcal{L}(t)$, $\mathtt{enumerate}(M)$ outputs (an STP equivalent to) $t$, or makes a recursive call $\mathtt{enumerate}(M^{\prime})$ for $M^{\prime}$ such that $M\subsetneq M^{\prime}\subseteq\mathcal{L}(t)$ (as we have already shown the termination of the algorithm).
By the completeness (Theorem 3.2) and minimality (Theorem 3.4), the value of $U$ initialized at the beginning of the call of $\mathtt{enumerate}(M)$ must contain $t^{\prime}$ such that $M\subseteq\mathcal{L}(t^{\prime})\subseteq\mathcal{L}(t)$. We perform case analysis on whether $t^{\prime}$ is picked up in the while-loop.
  • Suppose $t^{\prime}$ is eventually picked up in the while-loop. Then if $\mathcal{L}(t)=\mathcal{L}(t^{\prime})$, then $g_{\mathcal{F}}(t^{\prime})$ returns $\mathit{None}$ and $t^{\prime}$ is output. If $\mathcal{L}(t^{\prime})\subsetneq\mathcal{L}(t)$, then $g_{\mathcal{F}}(t^{\prime})$ returns Some(s~)Some($\widetilde${s}) such that $\widetilde{s}\in\mathcal{F}(\mathcal{L}(t^{\prime}))\subseteq\mathcal{F}(\mathcal{L}(t))\subseteq\mathcal{L}(t)$. Thus, $\mathtt{enumerate}(M^{\prime})$ is called for $M^{\prime}=M\cup\{\widetilde{s}\}\subseteq\mathcal{L}(t)$, as required.
  • Suppose $t^{\prime}$ is never picked up in the while-loop. Then $t^{\prime}$ must have been removed by the assignment $U:=U\setminus\{t^{\prime}\mid\widetilde{s}\in\mathcal{L}(t^{\prime})\}$. But then it must be the case that $\widetilde{s}\in\mathcal{L}(t^{\prime})$. Therefore, $\mathtt{enumerate}(M^{\prime})$ is called for $M^{\prime}=M\cup\{\widetilde{s}\}\subseteq\mathcal{L}(t^{\prime})\subseteq\mathcal{L}(t)$, as required.
A.1.9 Proofs for Section 3.4
Proof of Theorem 3.9.
@@
(1): By Lemma A.6, if $t\leadsto t^{\prime}$, then $t$ is solvable if and only if $t^{\prime}$ is solvable. By Lemma A.7, the length of a reduction sequence $t\leadsto t_{1}\leadsto\cdots\leadsto t_{n}$ is bounded by #t\#{t}. Thus, to check whether $t$ is solvable, it suffices to repeatedly reduce $t$ (in a non-deterministic manner) until no reduction is applicable, and check whether the resulting tuple pattern is of the form $(x_{1},\ldots,x_{k})$. As each step can be performed in polynomial time, the algorithm runs in polynomial time.
(3): Based on Lemma A.4, we can use the algorithm in Algorithm 1.
Function decide_inc($t_{1},t_{2}$):
 if $t_{2}$ is of the form $(x_{1},\ldots,x_{n})$ where $x_{1},\ldots,x_{n}$ are pairwise distinct then
    return True
   pick $t_{2}^{\prime}$ such that $t_{2}\leadsto t_{2}^{\prime}$ ;
 if $t_{1}^{\prime}:=\langle t_{2}\leadsto t_{2}^{\prime}\rangle(t_{1})$ is well defined then
    return decide_inc($t_{1}^{\prime},t_{2}^{\prime}$)
 else
    return False
Algorithm 1 Algorithm to decide $\mathcal{L}(t_{1})\stackrel{{\scriptstyle?}}{{\subseteq}}\mathcal{L}(t_{2})$.
@@
Here, we assume $|t_{1}|=|t_{2}|$; otherwise, we can immediately conclude that $\mathcal{L}(t_{1})\not\subseteq\mathcal{L}(t_{2})$. The algorithm repeatedly reduces $t_{2}$ and checks whether the reduction of $t_{2}$ can be “simulated by” the corresponding reduction of $t_{1}$. If $t_{2}$ is of the form $(x_{1},\ldots,x_{n})$, we can immediately conclude that $t_{1}\subseteq t_{2}$ holds, as $\mathcal{L}(t_{2})=\Sigma^{*}\times\cdots\times\Sigma^{*}$. Otherwise, we can always pick $t_{2}^{\prime}$ such that $t_{2}\leadsto t_{2}^{\prime}$ by the assumption that $t_{2}$ is solvable. By Lemma A.4, if $\langle t_{2}\leadsto t_{2}^{\prime}\rangle(t_{1})$ is well defined, then it suffices to check whether $\mathcal{L}(t_{1}^{\prime})\subseteq\mathcal{L}(t_{2}^{\prime})$; otherwise, we can immediately conclude that $\mathcal{L}(t_{1})\not\subseteq\mathcal{L}(t_{2})$ (on the last line).
@@
Each operation can be clearly performed in time polynomial in the size of the initial inputs $t_{1}$ and $t_{2}$. Since ${t_{2}}>\#{t_{2}^{\prime}}\geq 0$, the depth of the recursive call is bounded by $t_{2}$}. Thus, the algorithm runs in polynomial time. (As this proof shows, $t_{1}$ need not be solvable.)
@@
Before proving Theorem 3.10, we define the semantics of formulas. A valuation is a map from the set of variables to $\Sigma^{*}$. For a valuation $\rho$ and $w\in(\Sigma\cup\mathcal{V})^{*}$ such that $\mathbf{Vars}(w)\subseteq\mathit{dom}(\rho)$, we define $\rho w$ by:
$\rho\epsilon=\epsilon\qquad\rho(aw)=a\cdot(\rho w)\qquad\rho(xw)=\rho(x)\cdot(\rho w).$
For a valuation $\rho$ and a quantifier-free STP-formula $\varphi$, the relation $\rho\models\varphi$ is defined by induction on $\varphi$, as follows.